OpenStudy (anonymous):

The radius of a sphere is increasing at a rate of 3 inches per minute. a. Find the rate of change of the volume when r = 9 inches. b. find the rate of the circumference when r = 9 inches.

OpenStudy (anonymous):

AP Calc.

OpenStudy (anonymous):

\[V=4/3 \pi r^3\] take derivative w.r.t t on both sides: \[dV/dt= 4 \pi r^2 (dr/dt)\] when r=9 (dr/dt)=3 , dV/dt= 4*pi*81*3 for part b use same method on \[C=2\pi r\]

OpenStudy (anonymous):

You sir, are a life saver! I didn't get any of this stuff! Thanks so much!